Consider a constant total force of 1,216 N acting on an object. If the acceleration of the object is 9 m/s2, what is the mass in kg of this object if all other forces equal zero?
Answer #1

m = F /a = 135.11 Kg
Answer #2

F=ma

1216=9m

m=135.1111 kg
Answer #3

m = F/a = 1216 / 9 = 135.1 kg